# Hopf Bifurcation of Nonlinear Non-Hermitian Skin Effect

### v1

@@ II.2 Lattice model
|$\ket{\psi}=\ket{\psi^{0}}+\varepsilon\ket{\psi^{1}}+\varepsilon^{2}\ket{\psi^{2}}+\mathcal{O}\left(\varepsilon^{3}\right),\quad 1+\varepsilon a_{1}+\varepsilon^{2}a_{2}+\mathcal{O}\left(\varepsilon^{3}\right),$ (II.32)
@@
$\left(1+\varepsilon a_{1}+\varepsilon^{2}a_{2}\right)^{2}\left(\psi_{n+1}+\psi_{n-1}-2\psi_{n}\right)+2\varepsilon\left(1+\varepsilon a_{1}+\varepsilon^{2}a_{2}\right)\left(\left|\psi_{n}\right|^{2}-\tilde{\gamma}\right)\frac{\psi_{n+1}-\psi_{n-1}}{2}+2E\psi_{n}=0.$ (II.33)
@@
$\left(\psi_{n+1}^{0}+\psi_{n-1}^{0}-2\psi_{n}^{0}\right)+2E\psi_{n}^{0}=0,$ (II.34)
$\left(\psi_{n+1}^{1}+\psi_{n-1}^{1}-2\psi_{n}^{1}\right)+2E\psi_{n}^{1}=-2a_{1}\left(\psi_{n+1}^{0}+\psi_{n-1}^{0}-2\psi_{n}^{0}\right)-\left(\left|\psi_{n}^{0}\right|^{2}-\tilde{\gamma}\right)\left(\psi_{n+1}^{0}-\psi_{n-1}^{0}\right).$ (II.35)
The solution to Eq. (II.34) with the boundary condition $\psi_{0}^{0}=0$ is given as
$\psi_{n}^{0}=r_{0}\sin k_{0}n,\quad E\eqqcolon 1-\cos k_{0}\quad\left(r_{0}\in\mathbb{R},k_{0}>0\right).$ (II.36)
Substituting this into Eq. (II.35), we obtain
@@
$\quad=4a_{1}r_{0}\left(1-\cos k_{0}\right)\sin k_{0}n-2r_{0}\sin k_{0}\left[\left(\frac{r_{0}^{2}}{4}-\tilde{\gamma}\right)\cos k_{0}n-\frac{r_{0}^{2}}{4}\cos 3k_{0}n\right].$ (II.37)
@@
$a_{1}=0,\quad r_{0}=2\sqrt{\tilde{\gamma}},$ (II.38)
@@
$\psi_{n}^{1}=c\sin k_{0}n+\frac{\tilde{\gamma}^{3/2}}{\sin 2k_{0}}\left(\cos k_{0}n-\cos 3k_{0}n\right)\quad\left(c\in\mathbb{R}\right)$ (II.39)
@@
$\qquad=-2a_{2}\left(\psi_{n+1}^{0}+\psi_{n-1}^{0}-2\psi_{n}^{0}\right)-2\psi_{n}^{0}\psi_{n}^{1}\left(\psi_{n+1}^{0}-\psi_{n-1}^{0}\right)-\left(\left|\psi_{n}^{0}\right|^{2}-\tilde{\gamma}\right)\left(\psi_{n+1}^{1}-\psi_{n-1}^{1}\right).$ (II.40)
@@
$a_{2}=\frac{\tilde{\gamma}^{2}\left(3-4\cos^{2}k_{0}\right)}{8\cos k_{0}\left(1-\cos k_{0}\right)},\quad c=0.$ (II.41)
@@
$=2\sqrt{\tilde{\gamma}}\sin\left[k_{0}\left(1+\frac{3-4\left(1-E\right)^{2}}{8E\left(1-E\right)}\left(\tilde{\gamma}\varepsilon\right)^{2}\right)n\right]+\frac{\tilde{\gamma}^{3/2}\varepsilon}{\sin 2k_{0}}\left(\cos k_{0}n-\cos 3k_{0}n\right).$ (II.42)
@@
$k=k_{0}\left(1+\frac{3-4\left(1-E\right)^{2}}{8E\left(1-E\right)}\left(\tilde{\gamma}\varepsilon\right)^{2}\right)+\mathcal{O}\left(\varepsilon^{3}\right),$ (II.43)
@@
$E=1-\cos k-\frac{k\sin k\left(3-4\cos^{2}k\right)}{8\cos k\left(1-\cos k\right)}\left(\tilde{\gamma}\varepsilon\right)^{2}+\mathcal{O}\left(\varepsilon^{3}\right).$ (II.44)
While this expression recovers Eq. (II.31) for $\left|k\right|\ll 1$, it seems to break down for $k=\pm\pi/2$ as a consequence of lattice effects.
III Limit cycle for strong nonlinearity
We analyze a delocalized eigenstate associated with a limit cycle in phase space under the regime of strong nonlinearity $\varepsilon\gg 1$. Given that
$\partial_{x}^{2}\psi+2\left(\varepsilon\psi^{2}-\gamma\right)\partial_{x}\psi=\partial_{x}\left(\partial_{x}\psi+2\varepsilon\left(\frac{1}{3}\psi^{3}-\frac{\gamma}{\varepsilon}\psi\right)\right),$ (III.1)
the nonlinear Schrödinger equation for the nonlinear Hatano-Nelson model reads
$\partial_{x}\psi$ $=\varepsilon\left(\chi-2F\left(\psi\right)\right),$ (III.2)
$\partial_{x}\chi$ $=-\frac{2E}{\varepsilon}\psi,$ (III.3)
with
$F\left(\psi\right)\coloneqq\frac{1}{3}\psi^{3}-\frac{\gamma}{\varepsilon}\psi,\quad\chi\coloneqq\frac{1}{\varepsilon}\partial_{x}\psi+2F\left(\psi\right).$ (III.4)
Owing to $\varepsilon\gg 1$, Eq. (III.2) indicates that an arbitrary initial point in phase space $\left(\psi,\chi\right)$ rapidly converges to the nullcline $\chi=2F\left(\psi\right)$. Subsequently, it slowly drifts along this nullcline according to Eq. (III.3), ultimately leading to the convergence at the fixed point $\left(0,0\right)$ for $\gamma\leq 0$ and the emergence of a limit cycle for $\gamma>0$. Since the cubic function $\chi=2F\left(\psi\right)$ attains extrema ±$\pm\chi_{0}\coloneqq\pm\left(4/3\right)\left(\gamma/\varepsilon\right)^{3/2}$ at $\psi=\pm\psi_{0}=\pm\sqrt{\gamma/\varepsilon}$, the period of the limit cycle, corresponding to the wavelength of the delocalized eigenstate, is evaluated as
[…]
Defining a wavenumber as $k\coloneqq 2\pi/\lambda$, we have the energy dispersion relation
$E=\frac{3-2\log 2}{2\pi}\gamma k.$ (III.6)
Notably, despite the nonrelativistic nature of the original system, the strong nonlinearity gives rise to the emergent relativistic energy dispersion.
IV Nonlinear Hatano-Nelson model
IV.1 Critical nonlinear skin effect in the continuum
@@
$\partial_{x}^{2}\psi+2\left(\varepsilon\left|\psi\right|^{2}-\gamma\right)\partial_{x}\psi+2E\psi=0.$ (IV.1)
@@
IV.2 Nonlinear skin effect on the lattice
A discrete counterpart of Eq. (IV.1) is given as
$\left(\psi_{n+1}+\psi_{n-1}-2\psi_{n}\right)+2\left(\varepsilon\left|\psi_{n}\right|^{2}-\gamma\right)\frac{\psi_{n+1}-\psi_{n-1}}{2}+2E\psi_{n}=0,$ (IV.2)
@@ IV.2 Nonlinear skin effect on the lattice
$-\frac{1}{2}\left(1-\gamma+\varepsilon\left|\psi_{n}\right|^{2}\right)\psi_{n+1}-\frac{1}{2}\left(1+\gamma-\varepsilon\left|\psi_{n}\right|^{2}\right)\psi_{n-1}+\psi_{n}=E\psi_{n}.$ (IV.3)
@@
where |$\ket{\psi}=\left(\psi_{1},\cdots,\psi_{L}\right)^{T}$ denotes a single-particle eigenstate, and |$\ket{n}$ is a single-particle state localized at site $n$.
In Fig. S1 (b), we show the dependence of the localization length $\xi$ on the single-particle energy $E$. In the linear regime (i.e., $\varepsilon=0$), the localization length $\xi=1/\gamma$ of the skin state remains constant and exhibits no dependence on $E$. On the other hand, in the nonlinear regime, the system develops mobility edges at $E\simeq 0.7,1.3$; while $\xi$ is finite within the energy window $0.7\lesssim E\lesssim 1.3$, it diverges for $E\lesssim 0.7$, $E\gtrsim 1.3$. The presence of these mobility edges is reminiscent of those in the original (i.e., disordered) Hatano-Nelson model [112, *Hatano-Nelson-97].
Figure S1: (a) Wave function numerically obtained for the continuum nonlinear Hatano-Nelson model ($\gamma=0$, $\varepsilon=0.3$, $E=0.5$) with the boundary conditions $\psi\left(x=0\right)=0$, $\psi^{\prime}\left(x=0\right)=1$. Dotted line: $\psi\left(x\right)=3/\sqrt{x}$. (b) Inverse of the localization length, $1/\xi$, numerically obtained for the continuum nonlinear Hatano-Nelson model ($L=1000$, $\gamma=0.2$, $\varepsilon=0.3$) as a function of single-particle energy $E$.

### v2

@@ B.2 Lattice model
[…]
@@
$\left(1+\varepsilon a_{1}+\varepsilon^{2}a_{2}\right)^{2}\left(\psi_{n+1}+\psi_{n-1}-2\psi_{n}\right)+2\varepsilon\left(1+\varepsilon a_{1}+\varepsilon^{2}a_{2}\right)\left(\left|\psi_{n}\right|^{2}-\tilde{\gamma}\right)\frac{\psi_{n+1}-\psi_{n-1}}{2}+2E\psi_{n}=0.$ (B.18)
@@
$\left(\psi_{n+1}^{0}+\psi_{n-1}^{0}-2\psi_{n}^{0}\right)+2E\psi_{n}^{0}=0,$ (B.19)
$\left(\psi_{n+1}^{1}+\psi_{n-1}^{1}-2\psi_{n}^{1}\right)+2E\psi_{n}^{1}=-2a_{1}\left(\psi_{n+1}^{0}+\psi_{n-1}^{0}-2\psi_{n}^{0}\right)-\left(\left|\psi_{n}^{0}\right|^{2}-\tilde{\gamma}\right)\left(\psi_{n+1}^{0}-\psi_{n-1}^{0}\right).$ (B.20)
The solution to Eq. (B.19) with the boundary condition $\psi_{0}^{0}=0$ is given as
$\psi_{n}^{0}=r_{0}\sin k_{0}n,\quad E\eqqcolon 1-\cos k_{0}\quad\left(r_{0}\in\mathbb{R},k_{0}>0\right).$ (B.21)
Substituting this into Eq. (B.20), we obtain
@@
$\quad=4a_{1}r_{0}\left(1-\cos k_{0}\right)\sin k_{0}n-2r_{0}\sin k_{0}\left[\left(\frac{r_{0}^{2}}{4}-\tilde{\gamma}\right)\cos k_{0}n-\frac{r_{0}^{2}}{4}\cos 3k_{0}n\right].$ (B.22)
@@
$a_{1}=0,\quad r_{0}=2\sqrt{\tilde{\gamma}}$, (B.23)
@@
$\psi_{n}^{1}=c\sin k_{0}n+\frac{\tilde{\gamma}^{3/2}}{\sin 2k_{0}}\left(\cos k_{0}n-\cos 3k_{0}n\right)\quad\left(c\in\mathbb{R}\right)$ (B.24)
@@
$\qquad=-2a_{2}\left(\psi_{n+1}^{0}+\psi_{n-1}^{0}-2\psi_{n}^{0}\right)-2\psi_{n}^{0}\psi_{n}^{1}\left(\psi_{n+1}^{0}-\psi_{n-1}^{0}\right)-\left(\left|\psi_{n}^{0}\right|^{2}-\tilde{\gamma}\right)\left(\psi_{n+1}^{1}-\psi_{n-1}^{1}\right).$ (B.25)
@@
$a_{2}=\frac{\tilde{\gamma}^{2}\left(3-4\cos^{2}k_{0}\right)}{8\cos k_{0}\left(1-\cos k_{0}\right)},\quad c=0.$ (B.26)
@@
$=2\sqrt{\tilde{\gamma}}\sin\left[k_{0}\left(1+\frac{3-4\left(1-E\right)^{2}}{8E\left(1-E\right)}\left(\tilde{\gamma}\varepsilon\right)^{2}\right)n\right]+\frac{\tilde{\gamma}^{3/2}\varepsilon}{\sin 2k_{0}}\left(\cos k_{0}n-\cos 3k_{0}n\right).$ (B.27)
@@
$k=k_{0}\left(1+\frac{3-4\left(1-E\right)^{2}}{8E\left(1-E\right)}\left(\tilde{\gamma}\varepsilon\right)^{2}\right)+\mathcal{O}\left(\varepsilon^{3}\right),$ (B.28)
@@
$E=1-\cos k-\frac{k\sin k\left(3-4\cos^{2}k\right)}{8\cos k\left(1-\cos k\right)}\left(\tilde{\gamma}\varepsilon\right)^{2}+\mathcal{O}\left(\varepsilon^{3}\right).$ (B.29)
While this expression recovers Eq. (B.16) for $\left|k\right|\ll 1$, it seems to break down for $k=\pm\pi/2$ as a consequence of lattice effects.
Appendix C Nonlinear Hatano-Nelson model
C.1 Critical nonlinear skin effect in the continuum
@@
$\partial_{x}^{2}\psi+2\left(\varepsilon\left|\psi\right|^{2}-\gamma\right)\partial_{x}\psi+2E\psi=0.$ (C.1)
@@
C.2 Nonlinear skin effect on the lattice
A discrete counterpart of Eq. (C.1) is given as
$\left(\psi_{n+1}+\psi_{n-1}-2\psi_{n}\right)+2\left(\varepsilon\left|\psi_{n}\right|^{2}-\gamma\right)\frac{\psi_{n+1}-\psi_{n-1}}{2}+2E\psi_{n}=0,$ (C.2)
@@ C.2 Nonlinear skin effect on the lattice
$\frac{1}{2}\left(1-\gamma+\varepsilon\left|\psi_{n}\right|^{2}\right)\psi_{n+1}-\frac{1}{2}\left(1+\gamma-\varepsilon\left|\psi_{n}\right|^{2}\right)\psi_{n-1}+\psi_{n}=E\psi_{n}.$ (C.3)
@@
[…]
In Fig. S1 (b), we show the dependence of the localization length $\xi$ on the single-particle energy $E$. In the linear regime (i.e., $\varepsilon=0$), the localization length $\xi=1/\gamma$ of the skin state remains constant and exhibits no dependence on $E$. On the other hand, in the nonlinear regime, the system develops mobility edges at $E\simeq 0.7,1.3$; while $\xi$ is finite within the energy window $0.7\lesssim E\lesssim 1.3$, it diverges for $E\lesssim 0.7$, $E\gtrsim 1.3$. The presence of these mobility edges is reminiscent of those in the original (i.e., disordered) Hatano-Nelson model [115, *Hatano-Nelson-97].
Figure S1: (a) Wave function numerically obtained for the continuum nonlinear Hatano-Nelson model ($\gamma=0$, $\varepsilon=0.3$, $E=0.5$) with the boundary conditions $\psi\left(x=0\right)=0$, $\psi^{\prime}\left(x=0\right)=1$. Dotted line: $\psi\left(x\right)=3/\sqrt{x}$. (b) Inverse of the localization length, $1/\xi$, numerically obtained for the lattice nonlinear Hatano-Nelson model ($L=1000$, $\gamma=0.2$, $\varepsilon=0.3$) as a function of single-particle energy $E$.
Appendix D Nonlinear non-Hermitian Su-Schrieffer-Heeger model
As another prototypical example exhibiting the nonlinear skin effect, we consider a nonlinear non-Hermitian extension of the Su-Schrieffer-Heeger model [125]:
[…]
$H\left(k\right)=\left(v+w\cos k\right)\sigma_{x}+\left(w\sin k+\text{i}\gamma\right)\sigma_{y}=\begin{pmatrix}0&v+\gamma+we^{-\text{i}k}\\ v-\gamma+we^{\text{i}k}&0\end{pmatrix}$. (D.2)
[…]
$\left(v-\gamma+\varepsilon\,(\left|a_{n}\right|^{2}+\left|b_{n}\right|^{2})\right)b_{n}+wb_{n-1}$ $=Ea_{n},$ (D.3)
$\left(v+\gamma-\varepsilon\,(\left|a_{n}\right|^{2}+\left|b_{n}\right|^{2})\right)a_{n}+wa_{n+1}$ $=Eb_{n},$ (D.4)
which can be recast as
$\begin{pmatrix}a_{n+1}\\ b_{n+1}\end{pmatrix}=T\left(a_{n},b_{n}\right)\begin{pmatrix}a_{n}\\ b_{n}\end{pmatrix}$ (D.5)
with the nonlinear transfer matrix $T\left(a_{n},b_{n}\right)$ defined by
$T\left(a_{n},b_{n}\right)\coloneqq\begin{pmatrix}-\cfrac{v+\gamma_{\varepsilon}}{w}&\cfrac{E}{w}\\ -\cfrac{E\left(v+\gamma_{\varepsilon}\right)}{w\left(v-\gamma_{\varepsilon}\right)}&\cfrac{E^{2}-w^{2}}{w\left(v-\gamma_{\varepsilon}\right)}\end{pmatrix},\quad\gamma_{\varepsilon}\left(a_{n},b_{n}\right)\coloneqq\gamma-\varepsilon\left(\left|a_{n}\right|^{2}+\left|b_{n}\right|^{2}\right).$ (D.6)
The spatial distribution of the wave function for given energy $E$ is then determined via the successive application of the nonlinear transfer matrix $T\left(a_{n},b_{n}\right)$.
In the presence of nonlinearity $\varepsilon>0$, the wave function exhibits qualitatively distinct behavior depending on the sign of $\gamma$: it decays toward the right for $\gamma<0$ [Fig. S2 (a)], whereas it oscillates for $\gamma>0$ [Fig. S2 (b)]. This behavior is similar to that of the nonlinear Hatano-Nelson model and signals the Hopf bifurcation of the nonlinear skin effect. In Fig. S2 (c), we further show the localization length as a function of linear non-Hermiticity $\gamma$. In the linear regime, the localization length remains finite for arbitrary $\gamma\neq 0$, manifesting the skin effect. In the nonlinear regime, by contrast, the localization length diverges for $\gamma>0$, implying the delocalization due to the emergence of a limit cycle in the associated phase space.
Figure S2: Nonlinear non-Hermitian Su-Schrieffer-Heeger model ($v=1/3$, $w=1$, $E=1$). Boundary conditions: $a_{1}=b_{1}=1$. (a, b) Wave function for $\varepsilon=0.3$ with (a) $\gamma=-0.2$ and (b) $\gamma=0.2$. (c) Inverse of the localization length, $1/\xi$ ($\left|a_{n}\right|\sim e^{n/\xi}$), numerically obtained over $500\leq n\leq 1000$ $\varepsilon=0$ (gray), $\varepsilon=0.3$ (red)].
[…]